Can someone show me step by step in simple from.how to write a fraction or mixed number as a decimal. 5 7/50. 13/25
Answer by stanbon(75887)

Can someone show me step by step in simple from.how to write a fraction or mixed number as a decimal. 5 7/50. 13/25
----
In general you have to divide the numerator of the fraction part
by its denominator::
-----
But some conversions don't require that dividing::
5 7/50 = 5 14/100 = 5.14
---------
13/25 = (4*13)/(4*25 = 52/100 = 0.52
------------------
Cheers,
Stan H.
-------------